OpenStudy (anonymous):

Jordan has crayons. There are crayons in each box. One box is missing 2 crayons. How many boxes does she have?

OpenStudy (anonymous):

5

OpenStudy (anonymous):

Ten crayons in each box minus 2 crayons = 48 crayons 10b - 2 = 48 10b - 2 + 2 = 48 + 2 10b = 50 10b/10 = 50/10 b = 5 5 boxes.

OpenStudy (anonymous):

man thank you

OpenStudy (anonymous):

Yep!